3.4.14 13 cm with a standard deviation of 0.10 om For what lengths will a bolit be destroyed? Round to one decimal place as needed) OA A bolt will e destroyed f he length is less tham cmorgreater an cm O8. A boit will be destroyed t the legth is between cm and om OG. A bolt wll be destroyedit the legth is greater thanom A bolt will be destroyed i# the length is less than em
Answer #1

From the given data,

for 2 std deviation above value from mean =mean+2*std deviaiton =13+2*0.10=13.20 cm

2 std deviation below value from mean =mean-2*std deviaiton =13-2*0.10=12.80 cm

hence option A is correct. A bolt will be destroyed if the length is less than 12.80 or greater than 13.20 cm